OpenStudy (anonymous):

is there anyone who knows E=MC^2 elaboratly ? if so plz explain na..!

OpenStudy (anonymous):

oh..! i'm asking about some one in openstudy to explain about this...theory

OpenStudy (amistre64):

energy is directly proportional to mass. the c^2 is a constant of proportionality that makes it fit together :) or so ive been led to believe

OpenStudy (anonymous):

U want the derivation ???

OpenStudy (anonymous):

I know how to derive it ! and its principle

OpenStudy (anonymous):

i ll tomorrow.........

OpenStudy (anonymous):

I think that there is a constant added to a fuller version of the equation known as "y" in Greek lettering. It is known as the Lorentz factor and is basically the square root of 1 - (v/c)^2 or something similar. You can look up Lorentz factor.

OpenStudy (anonymous):

E = ymc^2 is an example used for obects closer to the speed of light. E = mc^2 is for objects moving at normal speeds.

OpenStudy (anonymous):

It relates the energy of a particle to its mass, and is a consequence of the theory of relativity. What it tells us is that mass \(m\) and energy \(E\) are equivalent, and that a little bit of mass is equal to a large amount of energy due to the multiplication factor of the speed of light c squared. It is this principle that permits power generation in nuclear power plants, the destructive energy of nuclear weapons, as well as the life giving energy of our sun. \(E=mc^2\) is only one part of the full equation which is far more interesting than that equation. This specific part only deals with particles at rest (the rest energy or rest mass), but there is also a part that deals with the object moving. The full equation is \[E=\sqrt{(pc)^2+(mc^2)^2}\] where \(p\) is the momentum of an object. There are three interesting things that comes from this equation. The first we have dealt with, is that of a particle at rest \(p=0\), meaning resulting in the realisation that energy and mass are equivalent. The second is that particles without mass, will still have a momentum, meaning that they will be able to move other particles about by colliding with them. Photons (particles of light) have zero rest mass, but are able to exert forces upon other objects through collision. Last but by far not the least, is that the equation predicts the anti-matter. the reason being comes from the mathematics. If you take the square root of any quantity, it will have two solutions. for example the square root of 4 i.e. \(\sqrt{4}\) will equal \(+2\) and \(-2\), because \((−2)(-2)=4\) as does \((+2)(+2)=4\). Therefore, the mathematics predicts that for a given momentum and rest mass, that there should exist particles with negative energy. Indeed, it was the Nobel prize winning physicist Paul Dirac who first took the second solutions seriously and predicted anti-matter a full year or two before it was experimentally observed.

OpenStudy (anonymous):

@jonnyMcA thnx....! gud one

OpenStudy (anonymous):

@TIMBUCK is that sooo? is that fr normal speed?

OpenStudy (anonymous):

@gita93 yu also plz derive it na....